### v2

# Local asymptotics of selection models with applications in Bayesian selective inference

Abstract
Contemporary focus on selective inference provokes interest in the asymptotic properties of selection models, as the working inferential models in the conditional approach to inference after selection. In this paper, we derive an asymptotic expansion of the local likelihood ratios of selection models. We show that under mild regularity conditions, they behave asymptotically like a sequence of Gaussian selection models. This generalizes the Local Asymptotic Normality framework of Le Cam (1960) to a class of non-regular models, and indicates a notion of local asymptotic selective normality as the appropriate simplifying theoretical framework for analysis of selective inference. Furthermore, we establish practical consequences for Bayesian selective inference. Specifically, we derive the asymptotic shape of Bayesian posterior distributions constructed from selection models, and show that they will typically be significantly miscalibrated in a frequentist sense, demonstrating that the familiar asymptotic equivalence between Bayesian and frequentist approaches does not hold under selection.
Keywords: Asymptotics; Local asymptotic normality; Frequentist calibration; Posterior distribution; Selective inference; Selection models.
1 Introduction
The techniques of selective inference aim to ensure validity in situations where the same data is used to select the inference to be considered and also to conduct it. Three main approaches have been proposed: the simultaneous approach (Berk et al., 2013), which ensures validity irrespective of the inference performed and the selection procedure; the ‘condition on selection’ approach (Fithian et al., 2017), where inference is conditioned on those hypothetical datasets which might have occurred which would have led to selection of the same inferential problem; and information splitting methods, such as data splitting, where information in the sample is split between that used for selection and that used for inference: see for example García Rasines and Young (2023).
In this work, we study the asymptotic behavior of selection models, which arise as the working inferential models under the conditional and information-splitting approaches to selective inference. These are statistical models of substantial contemporary practical importance in which the probability of observing or analyzing a sample depends on the sample itself. The term ‘selection model’ was coined by Fraser (1952, 1966), although their study dates back at least to Fisher (1934). Selection models are ubiquitous in statistical inference and have traditionally been used to model situations involving sampling bias, where samples are observed only if they satisfy a certain condition—for example, if they belong to a certain portion of the population (Rao, 1985). Selective inference uses a rather more general formulation, which is described in Section 2.
A significant body of methodological research in selective inference has been developed for Gaussian models, partly because they allow for tractable analytical procedures in some common contexts; see e.g. Lee et al. (2016). As a result, the theoretical properties of selection models within this setting are well understood, both from the frequentist and the Bayesian positions. The main contribution of this work is the identification of an asymptotic connection between general parametric selection models and Gaussian ones, thereby providing a uniform framework for analyzing the properties of the former and therefore deriving powerful theoretical insights.
More broadly, this paper contributes to the literature on non-standard asymptotics by establishing a limiting behavior which deviates from the standard Gaussian framework. We show that, under mild regularity conditions, a sequence of selection models from a parametric distribution is asymptotically equivalent to a sequence of selection models derived from a Gaussian location model, which, in contrast to standard asymptotics, can be highly asymmetrical if the selection effect is significant (Section 3). The relevant Gaussian models inherit a selection mechanism for the mean parameter that often depends on both the original model and the underlying data-generating parameter in an intricate manner, as illustrated in the examples of Sections 3.1 and 3.2. This showcases the unconventional nature of the analysis, compared to standard asymptotic theory, and underscores the need for careful consideration of the appropriate asymptotic framework for inference.
Our contribution can be regarded as a generalization of the Local Asymptotic Normality framework of Le Cam (1960) to the family of selection models. It indicates a notion of local asymptotic selective normality as the appropriate asymptotic framework for analysis of selective inference problems. In the context of selective inference, therefore, in some substantial generality, a parametric inference problem is asymptotically equivalent to an identifiable selective inference carried out on a location parameter under the Gaussian assumption, conducted from a single observation. As noted above, in contrast to the conventional Local Asymptotic Normality framework, the asymptotic inferential problem considered here can remain rather complicated due to the intricate nature of the relevant asymptotic selection mechanism.
As a natural corollary of the asymptotic approximation, we derive an extension of the Bernstein-von Mises Theorem to selection models (Section 4). We show that in a selection model with a fixed prior on the parameter, the Bayesian posterior is asymptotically equivalent to that obtained in a Gaussian selection model with a uniform prior on the location parameter. Crucially, this is a non-Gaussian limit, with significant practical consequences for selective inference, as the posterior in such a Gaussian selection model is miscalibrated from a frequentist perspective. In these models, frequentist calibration requires specification of a prior density which depends strongly on the sample size, as discussed in García Rasines and Young (2022).
2 Selection models
Let $Y^{n}=[Y_{1},\ldots,Y_{n}]$ be independent and identically distributed (IID) random variables from a distribution $F_{\theta}$ on a measurable space $(\mathcal{Y},\mathcal{A})$, where $\theta$ is a parameter from an open set $\Theta\subseteq\mathbb{R}^{p}$. The statistical model for the full sample, defined over $(\mathcal{Y}^{n},\mathcal{A}^{n})$, is $\mathcal{M}_{n}=\{F_{\theta}^{n}\colon\theta\in\Theta\}$. It is well known that, under mild regularity conditions, $\mathcal{M}_{n}$ can be asymptotically approximated by a Gaussian location model. Specifically, if $\{F_{\theta}\colon\theta\in\Theta\}$ is differentiable in quadratic mean (DQM) (van der Vaart, 1998, p. 93), then as $n\to\infty$ the sequence of local log-likelihood ratios admits the asymptotic expansion
$\log\prod_{i=1}^{n}\frac{f_{\theta+h_{n}/\sqrt{n}}(Y_{i})}{f_{\theta}(Y_{i})}=h^{T}I_{\theta}Z_{n}-\frac{1}{2}h^{T}I_{\theta}h+o_{F_{\theta}}(1)$
for every converging sequence $h_{n}\to h$, where $Z_{n}\rightsquigarrow N(0,I_{\theta}^{-1})$, $I_{\theta}$ is the Fisher information for a single sample, and $f_{\theta}(y)$ is the density of $F_{\theta}$ with respect to some fixed measure $\mu$ (we use $\rightsquigarrow$ to denote convergence in distribution). This property, formalized by Le Cam (1960), is known as Local Asymptotic Normality (LAN). Many key asymptotic results in parametric statistics can be derived from the LAN property, including the asymptotic distribution of maximum likelihood estimators and the Bernstein–von Mises theorem in Bayesian inference; see van der Vaart (1998) for a detailed account.
In wide generality, a selection model can be obtained from a base model $\mathcal{M}_{n}$ by conditioning the data on a random quantity $U_{n}$ such that $U_{n}\mid Y^{n}$ is distribution-constant. Typically, $U_{n}$ is either a function of $Y^{n}$, such as $\mathbf{1}(Y_{1}+\ldots+Y_{n}>0)$, where $\mathbf{1}(A)$ denotes the indicator function of the event $A$, or a randomized function of it, $u(Y^{n},W_{n})$, say, where $W_{n}$ is user-generated noise independent of $\theta$. For a realized value $U_{n}=u_{n}$, the conditional density of $Y^{n}\mid u_{n}$ is
$f_{\theta}(y^{n}\mid u_{n})=\frac{f(u_{n}\mid y^{n})}{f_{\theta}(u_{n})}\prod_{i=1}^{n}f_{\theta}(y_{i}),$ (1)
where $f(u_{n}\mid y^{n})$ and $f_{\theta}(u_{n})$ are the conditional and marginal densities of $U_{n}$ relative to some dominating measures. We denote the corresponding probability distribution by $F_{\theta}^{u_{n}}$. In our analysis, we will treat the sequence $\{u_{n}\colon n\in\mathbb{N}\}$ as fixed, as it represents the quantity being conditioned on in the analysis, so we will occasionally suppress $u_{n}$ from the notation. In particular, we will write
$f_{\theta}(y^{n}\mid u_{n})=\frac{p_{n}(y^{n})}{\varphi_{n}(\theta)}\prod_{i=1}^{n}f_{\theta}(y_{i}),$ (2)
where $p_{n}(y^{n})=f(u_{n}\mid y^{n})$ and $\varphi_{n}(\theta)=f_{\theta}(u_{n})$. This simplifies the expressions, in particular in Section 4 where we discuss Bayesian selective inference, and aligns better with the notation conventionally employed in most practical applications of selective inference.
Throughout this paper, the term selection model refers to any statistical model of the form $\{F_{\theta}^{u_{n}}:\theta\in\Theta\}$, derived from an underlying base model $\{F_{\theta}^{n}:\theta\in\Theta\}$ and a specified conditioning statistic $U_{n}$. Furthermore, by a Gaussian selection model is meant a selection model where the base model $F_{\theta}$ is Gaussian with a known variance matrix and an unknown mean $\theta\in\mathbb{R}^{p}$. In a generic selection model, we refer to $p_{n}(y^{n})$ as the selection function This encodes all relevant distributional information about the selection mechanism, in the sense that any two statistics and realized values $U_{n1}=u_{n1}$ and $U_{n2}=u_{n2}$ with proportional selection functions $p_{n1}(y^{n})\propto p_{n2}(y^{n})$ induce the same selection model on $Y^{n}$. We denote the corresponding sequence of selection models by $\mathcal{M}_{n}(u_{n})$.
Selection models play a central role in the study of sampling bias and selective inference. In the context of sampling bias, it is often assumed that the selection function factorizes into marginal selection functions for each observation, i.e. that $p_{n}(y^{n})=\prod_{i=1}^{n}p(y_{i})$, in which case the resulting selection model is IID and standard asymptotics apply under mild conditions. In selective inference, on the other hand, the selection condition typically acts on the dataset as a whole, and such a representation does not hold. Under the conditional approach to selective inference, if a subparameter $\psi=g(\theta)$ is only analyzed provided a pre-specified condition on the sample is satisfied, then inference on $\psi$ ought to be carried out conditionally on this event, commonly referred to as the selection event. Here, pre-specified means that the condition was decided by the statistician before analyzing the data. In the notation introduced above, the selection event takes the form $\{U_{n}=u_{n}\}$ for a specific value $u_{n}$, and its occurrence typically indicates that $\psi$ is worth investigating for a particular reason. For example, this could be rejection of some null hypothesis, indicating a significant effect size. We will see some common examples of selection events in the following section. It is worth noting that the conditioning operation in this case is not intrinsic to the data sampling mechanism, but a formal distributional correction undertaken to restore repeated-sampling validity to the inferences.
3 Asymptotic expansion of the likelihood
In this section, we establish a formal asymptotic connection between the sequence of selection models, $\{\mathcal{M}_{n}(u_{n})\colon n\geq 1\}$, and a sequence of Gaussian selection models. At the heart of the result is an asymptotic selection function that reflects how selection influences the model asymptotically. This can be regarded as providing an extension of the standard LAN framework, which can be recovered by considering a constant selection function, $p_{n}(y^{n})\propto 1$.
First, we examine how selection acts on Gaussian models. Let $Z\sim N(h,\Sigma)$, with parameter $h\in\mathbb{R}^{p}$ and fixed positive-definite covariance $\Sigma\in\mathbb{R}^{p\times p}$, and consider the selection model induced by an arbitrary selection function $p(z)$. The selective log-likelihood ratio for this model around $h=0$ is given by
$\log\frac{f_{h}(Z)}{f_{0}(Z)}=h^{T}\Sigma^{-1}Z-\frac{1}{2}h^{T}\Sigma^{-1}h+\log\frac{\varphi(0)}{\varphi(h)},\quad\varphi(h)=E_{h}[p(Z)].$ (3)
Note that the selective log-likelihood ratio follows itself a selective Gaussian distribution under any $h\in\mathbb{R}^{p}$, just as the log-likelihood ratio of a conventional, non-selective Gaussian model is itself Gaussian.
The main result of this paper, Theorem 1, establishes conditions under which the sequence of selective log-likelihood ratios of a given selection model $\mathcal{M}_{n}(u_{n})$ admits an asymptotic expansion of the form (3) for a certain specification of selection function, which in general depends on the true generating parameter $\theta$. It is worth remarking that, since selection functions can vary arbitrarily with $n$, $\{\mathcal{M}_{n}(u_{n})\colon n\in\mathbb{N}\}$ does not in general converge to a fixed Gaussian selection model. Instead, it can, under suitable conditions, be approximated by a sequence of such. Before stating the Theorem, we need to introduce some notation.
Notation. In the non-selective model $\mathcal{M}_{n}$, let $l_{\theta}(y_{i})=\log f_{\theta}(y_{i})$ denote the log-likelihood for the $i$-th observation, $\triangledown l_{\theta}(y_{i})$ the score, and $I_{\theta}=E_{\theta}[\triangledown l_{\theta}(Y_{i})\triangledown l_{\theta}(Y_{i})^{T}]$ the per-observation Fisher information, assumed to be positive-definite for all $\theta\in\Theta$. For a function $g:I\subseteq\mathbb{R}^{p}\to\mathbb{R}$, consider its Lipschitz norm,
$\|g\|_{\text{L}}=\sup\left\{\frac{|g(x)-g(y)|}{\|x-y\|}\colon x,y\in I,x\neq y\right\},$ (4)
[…]
$\|X-Y\|_{\text{BL}}=\sup\left\{E_{X}[g(X)]-E_{Y}[g(Y)]\colon\quad g\colon\mathbb{R}^{p}\to\mathbb{R},\quad\|g\|_{\text{BL}}\leq 1\right\}.$ (5)
For a probability distribution $\mu$ on $\mathbb{R}^{p},$ consider its Lebesgue decomposition: $\mu=\mu_{C}+\mu_{S}$, where $\mu_{C}$ is absolutely continuous with respect to the Lebesgue measure and $\mu_{S}$ is concentrated on a set of null Lebesgue measure. We say that $\mu$ has an absolutely continuous component if $\mu_{S}(\mathbb{R}^{p})<1$. Finally, the total variation distance between two probability distributions $\mu$ and $\nu$ on $\mathbb{R}^{p}$ is $\|\mu-\nu\|_{\text{TV}}=\sup\{|\mu(A)-\nu(A)|\colon A\in\mathcal{B}(\mathbb{R}^{p})\}$, where $\mathcal{B}(\mathbb{R}^{p})$ is the Borel sigma algebra on $\mathbb{R}^{p}$.
We are now equipped to state the main result of the paper.
Theorem 1.
Suppose that $\{F_{\theta}\colon\theta\in\Theta\}$ is DQM, let $Z_{n}\equiv Z_{n}(\theta)=n^{-1/2}I_{\theta}^{-1}\sum_{i=1}^{n}\triangledown l_{\theta}(Y_{i})$, and assume that the sequence of selection functions is uniformly bounded:
$\sup\{\|p_{n}\|_{\infty}\colon n\geq 1\}<\infty.$ (6)
For a fixed $\theta\in\Theta$, assume that $M(x)=E_{\theta}[\exp\{x^{T}\triangledown l_{\theta}(Y_{1})\}]<\infty$ in a neighborhood of the origin, that there exists a sequence of measurable functions $\{p^{*}_{n}(\cdot;\theta)\colon\mathbb{R}^{p}\to[0,\infty)\colon n\in\mathbb{N}\}$ satisfying $p^{*}_{n}(z;\theta)=E_{\theta}\left[p_{n}(Y^{n})\mid Z_{n}=z\right]$ for all $z$ in the support of $Z_{n}$, and that either of the following conditions are met:
  1. 1.
    The distribution of $\triangledown l_{\theta}(Y_{1})$ has an absolutely continuous component;
  2. 2.
    $\sup\{\|p^{*}_{n}(\cdot;\theta)\|_{\mathrm{BL}}\colon n\geq 1\}<\infty$.
Define
$\varphi^{*}_{n}(h)=E[p^{*}_{n}(Z;\theta)],\quad Z\sim N(h,I_{\theta}^{-1}).$ (7)
If $\inf\{\varphi_{n}(\theta)\colon n\geq 1\}>0$, then, for all convergent sequences $h_{n}\to h$,
$\log\frac{f_{\theta+h_{n}/\sqrt{n}}(Y^{n}\mid u_{n})}{f_{\theta}(Y^{n}\mid u_{n})}=h^{T}I_{\theta}Z_{n}-\frac{1}{2}h^{T}I_{\theta}h+\log\frac{\varphi_{n}^{*}(0)}{\varphi_{n}^{*}(h)}+o_{F^{u_{n}}_{\theta}}(1).$ (8)
Furthermore, if condition 1 is met, $\|Z_{n}-Z_{n}^{*}\|_{\mathrm{TV}}\to 0$ conditionally on $u_{n}$, where $Z_{n}^{*}$ follows a selective $N(0,I_{\theta}^{-1})$ distribution with selection function $p^{*}_{n}(z;\theta)$. Otherwise, $\|Z_{n}-Z_{n}^{*}\|_{\mathrm{BL}}\to 0$.
The proof of Theorem 1 can be found in the supplementary material along with the proofs of all the other theoretical results of the paper. It takes as starting point the LAN property of the underlying, non-selective model, from which the first two terms of the expansion can be derived. The main technical difficulty lies in validating the approximation of the probability ratios $\varphi_{n}(\theta+n^{-1/2}h)/\varphi_{n}(\theta)$ by their Gaussian counterparts. This requires consideration of the moment generating function of $Z_{n}$ given selection. The existence of $M(x)$, together with the regularity assumptions on the selection functions in the limiting Gaussian models, prove to be sufficient to ensure uniform validity. In addition, the conditions on the selection functions are fairly minimal, and satisfied in most standard applications of the framework, as demonstrated in the examples considered below.
The requirement that $\varphi_{n}(\theta)$ is bounded away from zero is needed to ensure that the Gaussian approximation of the normalized score remains valid after conditioning on selection and to preserve the scale of likelihood ratios. This condition guarantees asymptotic regularity of the model even if selection is deterministic, that is, when the selection functions are of the form $p_{n}(y^{n})=\mathbf{1}(y^{n}\in A_{n})$ for some events $A_{n}$. However, it is well-known that, in these scenarios, if $y^{n}$ falls close to the boundary of $A_{n}$, irregularities can arise, leading to poor inferential performance unless $n$ is very large.
Scenarios with $\varphi_{n}(\theta)\to 0$ can arise in models involving many parameters. While this asymptotic regime is not covered by the Theorem, in some scenarios it is possible to verify the boundedness condition either exactly or asymptotically:
  • (i)
    If the selection condition is rejection of a simple hypothesis $H_{0}\colon\theta=\theta_{0}$ at a level $\alpha$, then $\varphi_{n}(\theta)\geq\alpha$ for all $n$, provided the test is consistent. Thus, as long as $\alpha$ does vanish as $n\to\infty$, the condition is satisfied. This is typically the case for multiple comparisons models, such as the one described in Section 3.2, with a bounded number of parameters.
  • (ii)
    If the selection mechanism allows for the existence of a sequence of $\sqrt{n}$-consistent estimator $\hat{\theta}_{n}$ of $\theta$ conditionally on selection, then, for large values of $n$, the magnitude of $\varphi_{n}(\theta)$ can be assessed through that of $\varphi_{n}(\hat{\theta}_{n})$ (Lemma 1). This is always possible if selection acts on a subset of the observations, leaving the remaining ones free of selection bias (see Example 3.2 below), and often possible if it acts on a randomized version of the dataset (Examples 3.3 and 3.4). Examples of randomization mechanisms that enable estimators with the required accuracy are provided by Tian and Taylor (2018), Leiner et al. (2023), Neufeld et al. (2024), and Dharamshi et al. (2024).
Lemma 1.
If $\{F_{\theta}\colon\theta\in\Theta\}$ is DQM and there exists a sequence of estimators $\hat{\theta}_{n}$ such that $n^{1/2}\|\hat{\theta}_{n}-\theta\|=O_{F^{u_{n}}_{\theta}}(1)$, then $\varphi_{n}(\theta)=o(1)\iff\varphi_{n}(\hat{\theta}_{n})=o_{F^{u_{n}}_{\theta}}(1)$.
Selection rules that do not allow for the existence of $\sqrt{n}$-consistent estimators are typically deterministic and yield inferences with very limited power. Their lack of power often requires the collection of additional data to draw meaningful conclusions about the parameter, leading to a data carving scenario in which estimation of $\theta$ with the required accuracy is feasible.
It is also important to note that the selection function in the asymptotic model, $p^{*}_{n}(z;\theta)$, depends on the true parameter $\theta$. This dependency introduces additional complexity to the asymptotic analysis. Under the LAN framework, the true value of $\theta$ only influences the covariance of the limiting Gaussian location model, leaving the group structure unaltered and facilitating a unified approach to analyzing the properties of inferential procedures. By contrast, under selection, $\theta$ has a deeper structural influence on the limiting model through its presence in $p^{*}_{n}$. This may affect the asymptotic distribution of estimators and test functions, as well as optimal choice of inferential approach. It remains to be explored whether the true asymptotic selection function can be effectively approximated by $\hat{p}^{*}_{n}(z)=p^{*}_{n}(z;\hat{\theta})$, where $\hat{\theta}$ is a $\sqrt{n}$-consistent estimator of $\theta$.
The following examples are representative of many common situations encountered in selective inference.
Example 3.1.
[Deterministic selection] Let $Y_{1},\ldots,Y_{n}$ be an IID sample from an exponential family distribution with density
$f_{\theta}(y_{i})=h(y_{i})\exp\left\{\theta^{T}y_{i}-A(\theta)\right\},\quad\theta\in\Theta.$ (9)
Let $U_{n}=\mathbf{1}(\bar{Y}_{n}\in E_{n})$, where $\bar{Y}_{n}=n^{-1}(Y_{1}+\ldots+Y_{n})$ and $E_{n}\subseteq\mathbb{R}^{p}$ is an arbitrary sequence of Borel sets with a bounded number of connected components. For example, if $p=1$ and $E_{n}=[t_{n},\infty)$, such a selection condition corresponds to rejection of a null hypothesis about $\theta$ under a uniformly most powerful test. In this case, $Z_{n}=\sqrt{n}\triangledown^{2}A(\theta)^{-1}\{\bar{Y}_{n}-\triangledown A(\theta)\}$, and $p^{*}_{n}(z;\theta)=\mathbf{1}\{n^{-1/2}\triangledown^{2}A(\theta)z+\triangledown A(\theta)\in E_{n}\}$.
Example 3.2.
[Data carving] Consider the same setting as in the previous example, but with selection condition given by $\bar{Y}^{(1)}_{n}\in E_{n}$, where $\bar{Y}^{(1)}_{n}$ is the mean of the first $n/2$ samples, where $n$ is assumed to be even. Then,
$p^{*}_{n}(z;\theta)=\frac{1}{f_{Z_{n}}(z)}\int_{E_{n}}f_{\bar{Y}^{(1)}_{n}}\{2[n^{-1/2}\triangledown^{2}A(\theta)z+\triangledown A(\theta)]-\bar{y}_{1}\}f_{\bar{Y}^{(1)}_{n}}(\bar{y}_{1})\mathrm{d}\bar{y}_{1}.$ (10)
Condition 2 of the Theorem is satisfied provided $\sup\{\|\log f_{\bar{Y}^{(1)}_{n}}\|_{\text{L}}\colon n\geq 1\}<\infty$. Conditioning on selection based on a subsample of the data was introduced by Fithian et al. (2017) and is commonly referred to as data carving.
Example 3.3.
[Randomization] Consider the same setting, but suppose now that the selection mechanism acts on $U_{n}=\sqrt{n}\bar{Y}_{n}+W$, where the noise $W$ has a known density $f_{W}(w)$, so that selection takes the form $U_{n}\in E_{n}$. Then, $p^{*}_{n}(z;\theta)=P\{\sqrt{n}\bar{y}_{n}(z)+W\in E_{n}\}$, so
$\triangledown p^{*}_{n}(z;\theta)=\int_{E_{n}}-\triangledown^{2}A(\theta)\triangledown f_{W}\{u-\triangledown^{2}A(\theta)z-\sqrt{n}\triangledown A(\theta)\}\mathrm{d}u,$ (11)
which is bounded provided $f_{W}(w)$ is bounded. Proposed by Tian and Taylor (2018), randomized selection has become common in practice in post-selection inference, as it enables more powerful analyses.
Example 3.4.
[Condition on randomized statistic] Consider the same setting as in the previous example. In cases where knowledge of the precise form of $E_{n}$ is unavailable to the statistician, it is common to condition on the observed value of $U_{n}$ instead (García Rasines and Young, 2023), and the selection function becomes $p^{*}_{n}(z;\theta)=f_{W}\{u-\triangledown^{2}A(\theta)z+\sqrt{n}\triangledown A(\theta)\}$, which satisfies the required regularity conditions for most standard noise distributions, such as the standard Gaussian or Laplace distributions (Tian and Taylor, 2018).
3.1 A simple univariate example
Let us assume that the underlying model is exponential of rate parameter $\theta$, and that for scientific reasons we are interested in large values of $\theta$. That is, we would like to perform inference on $\theta$ only if the sample average, $\bar{Y}_{n}$, is small, indicating a large rate. We will illustrate the four previous types of selection mechanism in this context.
A deterministic selection mechanism would be of the form $\bar{Y}_{n}<t_{n}$ for some predetermined threshold value $t_{n}$, which might have been obtained through a power analysis or in some other data-independent way. For a given $\theta$ and realization $\bar{Y}_{n}=\bar{y}_{n}$, the normalized score for this model is $z=\theta^{2}\sqrt{n}(1/\theta-\bar{y}_{n})$, and the selection function in the asymptotic model becomes $p^{*}_{n}(z;\theta)=\mathbf{1}\{z>\theta^{2}\sqrt{n}(1/\theta-t_{n})\}$. If the analogous selection rule is applied only to half the samples, so that inference is performed whenever $\bar{Y}_{n}^{(1)}<t_{n}$, then, assuming $n$ is even and noticing that $Z_{n}$ follows a shifted and scaled gamma distribution, we obtain, after some manipulation,
$p^{*}_{n}(z;\theta)=\frac{\Gamma(n)}{\Gamma\left(\frac{n}{2}\right)^{2}}\left(\frac{\sqrt{n}\theta^{2}}{2\sqrt{n}\theta-2z}\right)^{n-1}\int_{0}^{t_{n}}\left[x\left(\frac{2}{\theta}-\frac{2z}{\sqrt{n}\theta^{2}}-x\right)\right]^{n/2-1}\mathrm{d}x.$
Finally, suppose selection acts on $U_{n}=\sqrt{n}\bar{Y}_{n}+W$, where $W\sim N(0,\sigma^{2}_{W})$ for some pre-determined $\sigma_{W}>0$, so that the sample gets selected if $U<t_{n}$. Then,
$p^{*}_{n}(z;\theta)=\Phi\left\{\frac{\sqrt{n}}{\sigma_{W}}\left(t_{n}-\frac{1}{\theta}\right)+\frac{z}{\sigma_{W}\theta^{2}}\right\},$
where $\Phi$ is the $N(0,1)$ distribution function. If, instead, we condition on the observed value of $U_{n}$, $u_{n}$ (perhaps due to unavailability of knowledge of the precise threshold $t_{n}$ applied), the selection function becomes
$p^{*}_{n}(z;\theta)=\phi\left\{\frac{1}{\sigma_{W}}\left(u_{n}-\frac{\sqrt{n}}{\theta}\right)+\frac{z}{\theta^{2}}\right\},$
with $\phi$ the $N(0,1)$ density function.
Figure 1 shows $r_{n}(h;\theta)=\log\{\varphi_{n}(\theta+h/\sqrt{n})/\varphi_{n}(\theta)\}$ and $r^{*}_{n}(h;\theta)=\log\{\varphi^{*}_{n}(h)/\varphi^{*}_{n}(0)\}$ in the four situations considered above, for $\theta=2$, $n=40$, and $h\in[-2,2]$. The main step in the proof of Theorem 1 is to establish that the two functions approximate each other uniformly in compact sets of $h$ as $n\to\infty$. This figure provides visual evidence of the approximation across the range of values of $h$ considered. In all cases, the functions $\varphi^{*}_{n}(h)$ where computed by numerical integration of $I_{\theta}^{1/2}\phi(I_{\theta}^{1/2}z)p^{*}_{n}(z;\theta)$ over $z$.
Figure 1: Blue: $r_{n}(h;\theta)$; black: $r_{n}^{*}(h;\theta)$. Top left: deterministic selection with $t_{n}=0.5$; top right: data carving with $t_{n}=0.5$; bottom left: randomization with $W\sim N(0,1)$ and $t_{n}=\sqrt{n}0.5$; bottom right: conditioning on $\sqrt{n}\bar{Y}_{n}+W=u_{n}$, with $W\sim N(0,1)$ and $u_{n}=3.8$.
3.2 Inference on winners
In this section we explore a classic example of selection bias, where inference is conducted on the most promising parameters chosen from a large set of candidates. This constitutes a paradigmatic application of selective inference in the IID setting. In such problems, $N$ independent populations under investigation, with respective parameters $\theta_{1},\ldots,\theta_{N}$, are ranked based on a data-dependent criterion (typically some measure of significance of the parameter), and only the top-ranked populations are selected for further analysis. For example, we may compute $N$ $p$-values of significance and select for formal inference only those parameters whose respective $p$-values are among the smallest $K<N$ or fall below a fixed significance threshold $t$. Once a subset of parameters has been selected, further data may be collected on the selected populations, leading to a data-carving scenario, as described in the previous section. This type of situation is common in clinical studies, where multiple treatments are tested and only the most promising ones are advanced for further investigation. Another common application occurs in genomics, where a vast number of gene expressions are measured, and only those showing the strongest signals, such as significant associations with a disease, are formally analyzed.
We will consider the following specification of the problem. For $j\in\{1,\ldots,N\}$ and $i\in\{1,\ldots,n_{1}\}$, let $Y_{i}^{j}\sim N(\mu_{j},\sigma^{2}_{j})$ independently, where all $\mu_{j}$’s and $\sigma^{2}_{j}$’s are unknown. That is, we have random samples of size $n_{1}$ from $N$ different Gaussian populations. Suppose that we are only interested in means with large values. Accordingly, we compute $N$ $t$-statistics, $\{S_{j}=\bar{Y}^{j}/\sqrt{V^{j}/n_{1}}\colon j=1,\ldots,N\}$, where $\bar{Y}^{j}$ and $V^{j}$ are the maximum likelihood estimators of $\mu_{j}$ and $\sigma_{j}^{2}$, and keep the populations with the largest statistics for inference. We will not analyze any specific choice of selection criterion because, as we shall see, all of them are operationally identical under the considered form of inference. Thus, without loss of generality, suppose that $\mu_{j}$ is selected for inference if $S_{j}>T_{n_{1},N}(S_{-j})\equiv T_{j}$, where $S_{-j}$ contains all $t$-statistics except $S_{j}$. This threshold could be $T_{j}=\max(S_{-j})$, or it could be a constant significance threshold obtained from a formal hypothesis test, possibly with a multiplicity correction if $N$ is large. Furthermore, assume that, from each of the selected populations, we collect $n_{2}$ extra samples to increase inferential power: $Y_{n_{1}+1}^{j},\ldots,Y_{n}^{j}$, where $n=n_{1}+n_{2}$ is the total sample size. This is not necessary for the discussion, but mimics standard procedure in many applications.
In problems of this type, it is common to conduct inference on each of the selected means conditionally on the data from the other $N-1$ populations, in addition to conditioning on the selection event, so as to eliminate the $2(N-1)$ nuisance parameters. For example, let us assume that $\mu_{1}$ is one of the selected means. Selective inference on $\mu_{1}$ is then to be conducted from the model $Y_{1}^{1},\ldots,Y_{n}^{1}\sim N(\mu_{1},\sigma_{1}^{2})$ with selection event $S_{1}>t_{1}=T_{1}(s_{2},\ldots,s_{N})$, which only depends on the remaining populations through the (now constant) threshold $t_{1}$. In a selective many-parameter problem, such reduction is often possible and enables analysis of the high-dimensional model in terms of a few independent problems involving low-dimensional selected parameters. Therefore, in our empirical investigations, we shall simply consider selection models of the form $Y_{1},\ldots,Y_{n}\sim N(\mu,\sigma^{2})$ with selection event $S>t$, for some fixed $t\in\mathbb{R}$, where the selected population index has been dropped. The parameter is $\theta=(\mu,\sigma^{2})$.
Let $(\bar{Y}_{n},V_{n})$ be the maximum likelihood estimator of $\theta$. We have
$Z_{n}=(Z_{n}^{1},Z_{n}^{2})\rightsquigarrow N\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}\sigma^{2}&0\\ 0&2\sigma^{4}\end{pmatrix}\right),$
where $Z_{n}^{1}=\sqrt{n}(\bar{Y}_{n}-\mu)$ and $Z_{n}^{2}=\sqrt{n}(V_{n}-\sigma^{2}+(\bar{Y}_{n}-\mu)^{2})$. The selection function is $p_{n}(y^{n})=\mathbf{1}(s>t)$, where $s$ is the $t$-value as defined above (using only the first $n_{1}$ observations). The selection function of the asymptotic Gaussian model on $Z_{n}$ is $p_{n}^{*}(z;\theta)=\mathbb{E}_{\theta}[p_{n}(Y^{n})\mid Z_{n}=z]$, which does not admit, to the best of our knowledge, a simple analytical expression. Despite the apparent simplicity of the model, derivation of the appropriate asymptotic model for inference is complex, as the asymptotic selection function is the conditional acceptance probability of a subsample $t$-test given the full-sample sufficient statistics $(\bar{Y}_{n},V_{n})$. This, in turn, induces a selection probability function $\varphi_{n}^{*}(h)$ on a Gaussian location model $Z\sim N(h,I_{\theta}^{-1})$ which is non-standard: it cannot be represented by a standard selection criterion on the Gaussian model, such as involving a significance test on $h$. Figure 2 shows one-dimensional cuts of $p^{*}_{n}(z;\theta)$, $r_{n}(h;\theta)=\log\{\varphi_{n}(\theta+h/\sqrt{n})/\varphi_{n}(\theta)\}$ and $r_{n}^{*}(h;\theta)=\log\{\varphi^{*}_{n}(h)/\varphi^{*}_{n}(0)\}$ for $n_{1}=20$, $n_{2}=20$, $\theta=(0,1)$ and $t=1$. Under the true parameter, the selection probability is $\varphi_{n}(0,1)\approx 0.16$. The first two functions were computed via numerical integration, and $r_{n}^{*}(h;\theta)$ via Monte Carlo integration of $p^{*}_{n}(z;\theta)$. As expected, the approximation of the true selection probability by a Gaussian probability is much more accurate in the direction of the mean ($h_{1}$) than in the direction of the variance ($h_{2}$): in this problem $Z_{n}^{1}$ is exactly normally distributed. Additionally, since the selection condition is employed to determine large values of $\mu$, the selection function fluctuates much more in the first coordinate, corresponding to $\mu$ (top left plot), than in the second one (top right).
Figure 2: Top left: $p^{*}_{n}(z_{1},0;\theta)$; top right: $p^{*}_{n}(1,z_{2};\theta)$; bottom left: $r_{n}(h_{1},0;\theta)$ (blue) and $r^{*}_{n}(h_{1},0;\theta)$ (black); bottom right: $r_{n}(1,h_{2};\theta)$ (blue) and $r^{*}_{n}(1,h_{2};\theta)$ (black), with $\theta=(0,1)$.
4 Asymptotic posterior distributions
Equipped with the formal expansion provided by Theorem 1, we are now in a position to study the asymptotic shape of posterior distributions constructed from selection models. To this end, suppose we have a prior distribution $\pi(\theta)$ over $\Theta$. Upon observing a sample $Y^{n}=y^{n}$ from a selection model $\mathcal{M}_{n}(u_{n})$, the posterior density is given by
$\pi_{n}(\theta\mid y^{n})=\frac{\pi(\theta)}{c_{n}(y^{n})}f_{\theta}(y^{n}\mid u_{n})\propto\frac{\pi(\theta)}{\varphi_{n}(\theta)}\prod_{i=1}^{n}f_{\theta}(y_{i}),$ (12)
where $c_{n}(y^{n})$ is the marginal density of $Y^{n}$ in the selection model. We deliberately avoid the notation $\pi_{n}(\theta\mid y^{n},u_{n})$, as it leads to confusion with the standard, non-selective posterior
$\pi_{n}(\theta\mid y^{n},u_{n})\propto\pi(\theta\mid u_{n})f_{\theta}(y^{n}\mid u_{n})\propto\pi(\theta)f_{\theta}(y^{n}).$ (13)
Below we elaborate further on the distinction between these posteriors.
Throughout this section, we adopt the frequentist stance that there is a true underlying data-generating parameter, which is denoted $\theta_{0}$ so as to avoid confusion with other values of $\theta$ at which the posterior density is evaluated. Additionally, for a true parameter value $\theta_{0}\in\Theta$, we denote by $\pi_{h,n}(h\mid y^{n})$ the posterior density for the local parameter $h=\sqrt{n}(\theta-\theta_{0})$. The corresponding posterior distribution is
$\Pi_{h,n}(B\mid y^{n})=\int_{B}\pi_{h,n}(h\mid y^{n})\mathrm{d}h,\quad B\in\mathcal{B}(\mathbb{R}^{p}).$ (14)
Posterior distributions of this form are sometimes called selective posteriors or selection-adjusted posteriors. By contrast to the regular, non-selective IID case, these distributions have a factor of $1/\varphi_{n}(\theta)$ in their density which, as we shall see, can significantly modify their behavior. Selection-adjusted posterior distributions have appeared in multiple contexts; see García Rasines and Young (2022) for a detailed discussion. In selective inference, they arise under the so-called ‘fixed parameter’ regimes, as introduced by Mandel and Rinott (2007, 2009) and further developed by Yekutieli (2012). These are notional joint sampling regimes of $(\theta,Y)$ where $\theta$ is sampled from the prior, fixed, and data is generated conditionally on $\theta$ until the selection condition ($U_{n}=u_{n}$) is satisfied. Under such sampling regimes it is argued that the posterior density ought to be constructed by combining the prior $\pi(\theta)$ with the likelihood of the conditional model $\mathcal{M}_{n}(u_{n})$, yielding (12). By contrast, ‘random parameter’ regimes arise when $(\theta,Y)$ are sampled jointly until selection occurs. In the latter case, the usual posterior density $\pi_{n}(\theta\mid y^{n})\propto\pi(\theta)\prod_{i=1}^{n}f_{\theta}(y_{i})$, blind to selection, is recommended instead for inference (Yekutieli, 2012). An interesting conceptual discussion is also provided by Harville (2021), who notes that (12) follows from a standard Bayesian updating of the modified prior $\pi^{*}(\theta)=\pi(\theta)/\varphi(\theta)$. This has a natural interpretation in the selective inference context: Bayesian inference based on a selection-unadjusted prior has very poor frequentist behavior for $\theta$’s where $\varphi(\theta)$ is small. Thus, by increasing the prior density at those values, this selection-corrected prior provides some level of protection against selection effects.
Selection-adjusted posteriors are rapidly gaining popularity in theory and practice. Within the regression framework, their utility is highlighted in Panigrahi and Taylor (2018), Panigrahi et al. (2020) and Panigrahi et al. (2021). Objective Bayes procedures based on this type of posterior construction are discussed in Woody et al. (2021). Recent real-data applications include MacKinnon and Pavlovic (2022), who employed them to analyze hop market prices, and Panigrahi et al. (2023), who employed them to uncover important gene pathways in a radiogenomic analysis.
The growing popularity of selection-adjusted posteriors underscores the importance of understanding their theoretical properties, particularly as their behaviour deviates from that of standard Bayesian models. This section makes two primary contributions. We prove that the consequences of Theorem 1 extend to the Bayesian framework, leading to a result analogous to the Bernstein von-Mises Theorem for the context of selection. Specifically, we show that the asymptotic shape of the posterior distribution matches that of the corresponding Gaussian selection model under a uniform prior. These latter models provide a more intuitive framework that can be leveraged to better understand the former. Furthermore, as a consequence of the previous result, we indicate a key practical limitation of Bayesian selective inference: probabilistic claims about the parameter (e.g. that it lies within a certain credible interval with probability 90%), cannot be given a frequentist interpretation with high accuracy. In order to achieve frequentist-matching properties, more complex prior distributions that depend on the sample size are required.
4.1 Asymptotic behavior of selective posteriors
By Theorem 1, under certain conditions, selection models behave asymptotically as selective Gaussian models with observation $Z\sim N(h,I_{\theta_{0}}^{-1})$ and selection function $p_{n}^{*}(z;\theta_{0})$. It is thus natural to expect that under a similar set of assumptions the corresponding posterior distributions match asymptotically. To formalize this equivalence, we define the posterior density under the latter model with the improper uniform prior $\pi(h)\propto 1$ as
$\pi^{*}_{n}(h\mid z)=\frac{\text{det}(I_{\theta_{0}})^{1/2}}{\varphi^{*}_{n}(h)c^{*}_{n}(z)}\phi\left\{\|I_{\theta_{0}}^{1/2}(h-z)\|\right\},$ (15)
where $c^{*}_{n}(z)$ is a normalizing constant, and denote by $\Pi^{*}_{n}(\cdot\mid z)$ the corresponding probability distribution. The following is a direct generalization of Theorem 10.1 in van der Vaart (1998), and the proof follows largely the same steps. First, we show that the approximation holds conditionally on a sequence of balls of decreasing radius around $\theta_{0}$, which follows from Theorem 1. Then, we show that the posterior probability outside these balls is asymptotically negligible, which is guaranteed by the lower-bound assumption on $\varphi_{n}(\theta_{0})$.
Proposition 1.
Consider a sequence of selection models and a true parameter $\theta_{0}\in\Theta$ satisfying the assumptions of Theorem 1, and assume that the prior distribution is absolutely continuous, with a density $\pi(\theta)$ which is continuous and positive at $\theta_{0}$. Furthermore, suppose that there exists a sequence of tests $T_{n}\colon\mathcal{Y}^{n}\to[0,1]$ such that, for all $\varepsilon>0$,
$E_{\theta_{0}}(T_{n})\to 0\quad\text{and}\quad\sup_{\|\theta-\theta_{0}\|\geq\varepsilon}E_{\theta_{0}}(1-T_{n})\to 0.$ (16)
Then, the sequence of posterior distributions satisfies
$\left\|\Pi_{h,n}(\cdot\mid Y^{n})-\Pi^{*}_{n}(\cdot\mid Z_{n})\right\|_{\mathrm{TV}}\xrightarrow{F_{\theta_{0}}^{u_{n}}}0,$ (17)
where $Z_{n}=n^{-1/2}I_{\theta}^{-1}\sum_{i=1}^{n}\triangledown l_{\theta}(Y_{i})\rvert_{\theta_{0}}$. That is, the Total Variation distance between the true and asymptotic posteriors vanishes asymptotically in probability conditionally on selection.
When the data-generating parameter is $\theta_{0}$, Theorem 1 indicates that $Z_{n}$ follows asymptotically a selective Gaussian distribution with selection function $p^{*}_{n}(z;\theta_{0})$. Hence, Proposition 1 provides a framework for understanding the frequentist behavior of selective Bayesian methods by examination of their asymptotic, and simpler, Gaussian equivalents. It is important to note that the limiting posterior distributions $\Pi^{*}_{n}(\cdot\mid Z_{n})$ are not Gaussian unless $p^{*}_{n}(z;\theta)$, and thus $\varphi_{n}^{*}(h)$, are constant, as selection modifies the likelihood through the factor $1/\varphi_{n}^{*}(h)$. The impact of this factor is often manifested in the form of higher posterior variance and heavier tails than those observed in the absence of selection, and, in some cases, pronounced skewness. The latter effect is particularly notable when $\varphi_{n}^{*}(h)$ is strictly monotone, reflecting a selection mechanism that systematically favors either large or small values of the location parameter. Additionally, the impact of selection on the posterior grows with the amount of information used for selection. Consequently, hard-truncation mechanisms, such as the ones illustrated in Example 3.1, which employ all available data for selection, produce stronger effects than carved or randomized mechanisms (Examples 3.2 and 3.3), as can be observed in Figure 3.
Figure 3 compares the posterior densities for the local parameter $h=\sqrt{n}(\theta-\theta_{0})$, $\pi_{h,n}(h\mid y^{n})$, for two exponential models with their theoretical Gaussian limits. We examine two of the scenarios described at the end of Section 3: deterministic selection and randomization with Gaussian noise. The model is $Y_{1},\ldots,Y_{50}\sim\text{Exp}(\theta)$, $\theta\sim\text{Gamma}(1,0.1)$, and the true parameter is $\theta_{0}=2$. All posterior densities correspond to an observed sample average of $\bar{y}_{n}=0.45$. The true posterior densities based on the exponential likelihood, $\pi_{h,n}(h\mid y^{n})$, are shown in blue, while the Gaussian approximations $\pi^{*}_{h,n}(h\mid z_{n})$, as predicted by the theorem, are shown in black. For comparison, the posterior density in the absence of selection, which is approximately a rescaled $N(1/\bar{y}_{n},\theta_{0}^{2}/n)$ under standard theory, is shown in red. Two selection mechanisms are considered: a deterministic one (left), with selection condition $\bar{Y}_{n}<0.5$, and a randomized one (right), with selection condition $\bar{Y}_{n}+0.2\times N(0,1/n)<0.5$. In both cases, the selection probabilities $\varphi_{n}(\theta)$ and $\varphi^{*}_{n}(h)$ were obtained via numerical integration, as in the numerical investigation presented in Figure 1. We can readily observe that both posterior densities are non-Gaussian, exhibit higher uncertainty than the non-selective ones, have slowly decaying tails in the direction of the parameter space where $\varphi^{*}_{n}(\theta)$ vanishes, and demonstrate marked skewness, particularly in the case involving deterministic selection.
Figure 3: Realizations of the posterior densities in the exponential model (blue), their Gaussian approximations (black), and the standard, non-selective posteriors (red). Left: deterministic selection; right, randomized selection.
For further numerical illustration, we revisit the example discussed in Section 3.2, inference on a normal mean, under a selection condition defined by a $t-$ test on a subsample. We assume true parameter value $\theta_{0}=(0,1)$ and consider sample sizes $n_{1}=100,n_{2}=50$, with threshold $t_{n}=1$, which corresponds to a selection probability of about 0.15. Figure 4 illustrates, for a particular data sample, the marginal posteriors for $h_{1}$ and $h_{2}$, derived from the exact posterior $\pi_{h,n}(h\mid y^{n})$, constructed assuming the prior $\pi(\mu,\sigma^{2})\propto 1/\sigma$, and the Gaussian approximation $\pi^{*}_{h,n}(h\mid z_{n})$. In each case, the full posterior for the local parameter $h=(h_{1},h_{2})$ is constructed by MCMC sampling. The vertical lines in the figure indicate the limits of 90% credible intervals for $h_{1},h_{2}$ constructed from the exact posterior $\pi_{h,n}(h\mid y^{n})$. In the case of $h_{1}$, for instance, this interval is $(-2.11,2.50)$ and has probability content under the marginal posterior corresponding to the Gaussian approximation of 0.905. The corresponding figure in the case of $h_{2}$ is 0.884. The marginal posteriors in the true selection model and the Gaussian approximation model match closely, but not exactly.
Figure 4: Realizations of the exact marginal posterior densities of $h_{1}$ and $h_{2}$ in the inference on winners model (blue) and their Gaussian approximations (black). Exact marginal 90% credible intervals shown by vertical lines.
We repeated the same analysis for a series of 1000 replications, for several combinations of sample sizes $(n_{1},n_{2})$ and selection threshold $t_{n}$. Table 1 averages the probability content of the exact 90% credible interval under the approximate Gaussian posterior over the replications, with the standard deviations over the replications of this given in parentheses. Though the simulation is limited, there are some broad conclusions that can be drawn. The probability content averages are close to 90%, even for small sample sizes, though there is evidence that a larger sample size is required for matching of the posterior distributions for $h_{2}$. The probability content figures are more stable over replications for the parameter $h_{2}$ than the parameter $h_{1}$, where variability is seen to increase as the selection threshold $t_{n}$ increases, so that the selection probability shrinks. This is to be expected, as selection primarily affects the mean $\mu$, favouring samples which provide evidence of larger $\mu$, and therefore has greater impact on inference for the local mean $h_{1}$ than the nuisance parameter $h_{2}$.
Table 1: Average probability content of marginal 90% credible intervals, over 1000 replications.
$(n_{1},n_{2})$ $t_{n}$ $h_{1}$ $h_{2}$
$(20,20)$ 0 0.898 (0.051) 0.850 (0.075)
1 0.905 (0.047) 0.869 (0.063)
2 0.916 (0.039) 0.906 (0.054)
3 0.914 (0.048) 0.924 (0.064)
$(50,25)$ 0 0.907 (0.036) 0.871 (0.055)
1 0.917 (0.031) 0.888 (0.047)
2 0.913 (0.054) 0.903 (0.048)
3 0.892 (0.099) 0.917 (0.051)
$(100,50)$ 0 0.908 (0.027) 0.884 (0.040)
1 0.915 (0.029) 0.892 (0.037)
2 0.918 (0.045) 0.905 (0.034)
3 0.886 (0.130) 0.909 (0.039)
4.2 Frequentist calibration of selection-adjusted Bayesian inference
In the classical IID framework, the Bernstein-von Mises Theorem establishes asymptotic equivalence between the Bayesian and frequentist approaches to inference on $\theta$. In practical terms, this implies that Bayesian posterior probabilities can be interpreted as frequentist probabilities under the assumption that there is a fixed, true data-generating parameter $\theta_{0}$. For instance, $1-\alpha$ Bayesian credible intervals are also asymptotic $1-\alpha$ frequentist confidence intervals. This equivalence stems from the fact that Gaussian location models with a uniform prior satisfy this probability-matching condition exactly.
In the presence of selection, this is no longer the case: selection-adjusted posterior distributions can remain significantly miscalibrated as the sample size increases. This follows by combining Proposition 1 with the observation that the uniform prior on the location parameter lacks a Bayesian-frequentist equivalence under selection (Proposition 2). Achieving approximate frequentist calibration in this setting typically requires the use of more carefully tuned priors that account for the selection mechanism and vary with the sample size: see García Rasines and Young (2022). This is not surprising, since, typically in selective inference, the dependence of the selection function on $n$ does not vanish asymptotically. This consideration of mis-calibration is relevant both for Bayesians and for frequentist practitioners who use Bayesian methods for their computational or interpretive convenience.
In general, given a one-dimensional parameter of interest $\psi=g(\theta)$, a prior density for $\theta$ is probability-matching if $\Pi(g(\theta_{0})\mid Y)\sim U(0,1)$ under repeated sampling of $Y$ when $\theta_{0}$ is the true parameter, where $\Pi(\cdot\mid Y)$ is the marginal posterior distribution function of $\psi$ induced by the said prior. This ensures that credible intervals are valid confidence intervals, as
$P_{\theta_{0}}\{\Pi^{-1}(\alpha_{1}\mid Y)\leq\psi_{0}\leq\Pi^{-1}(\alpha_{2}\mid Y)\}=P_{\theta_{0}}\{\alpha_{1}\leq\Pi(\psi_{0}\mid Y)\leq\alpha_{2}\}=\alpha_{2}-\alpha_{1},$
for all $\theta_{0}$ and $10<\alpha_{1}<\alpha_{2}<$ 1. Conveniently, improper uniform priors on location parameters are probability-matching. Thus, by the Bernstein-von Mises Theorem, any fixed prior density which is strictly positive at $\theta_{0}$ is first-order probability matching, meaning that the distribution of $\Pi(\psi_{0}\mid Y)$ converges to a $U(0,1)$ at a rate of $1/\sqrt{n}$, and hence credible intervals are asymptotically valid confidence regions. The following proposition shows that, in wide generality, in a selective Gaussian model, unlike the non-selective counterpart, there is no fixed prior which is, even asymptotically, probability matching. This includes the uniform prior $\pi(\theta)\propto 1$. Consequently, the limiting Gaussian posteriors indicated in Corollary 1 are not calibrated in a frequentist sense.
Proposition 2.
Let $Y\sim N(\theta,\Sigma)\in\mathbb{R}^{p}$, with $\theta\in\mathbb{R}^{p}$ unknown and $\Sigma$ known and positive definite. Let $p(y)$ be a selection function such that
$\sum_{j=1}^{p}\text{Cov}(Y_{i},Y_{j})\frac{\partial}{\partial\theta_{j}}\varphi(\theta)\neq 0\quad\text{for all }\theta\in\mathbb{R}^{p}.$
Then, no prior density is probability-matching for $\theta_{i}$ in the selective model for $Y$.
This proposition relates to a well-known result due to Lindley (1958), which states that the only exact probability-matching priors in one-dimensional models are uniform priors for location parameters. In this case, the condition assumed that $\varphi(\theta)$ is not constant along the lines $\{[\text{Cov}(Y_{i},Y_{1}),\ldots,\text{Cov}(Y_{i},Y_{p})]t\colon t\in\mathbb{R}\}$ prevents the selective model from being a location model in these directions. The proof proceeds by finding a transformation of the original parameter $\theta\to(\theta_{i},\chi)$ such that $\theta_{i}$ and $\chi$ are orthogonal, and imposing that $\varphi$ not be constant in $\theta_{i}$ in this parametrization. Although seemingly complex, this condition ought to be satisfied by any reasonable selection rule which is used to identify interesting values of the parameter of interest $\theta_{i}$. If $Y_{i}$ is uncorrelated with the remaining variables, the required condition is simply
$\frac{\partial}{\partial\theta_{i}}\varphi(\theta)\neq 0.$
The degree of departure from uniformity of the marginal posterior density will generally depend on how strong selection bias is, that is how much the selection mechanism alters the distribution of the data under the true parameter, measured through the partial derivatives of $\varphi(\theta)$.
The lack of frequentist calibration can be made more precise in the one-dimensional setting, when the selection function of the Gaussian model is monotonic. In such cases, the posterior distribution tends to overstate, under repeated sampling, regions of the parameter space with low selection probability, i.e. it overcompensates selection bias. The following result extends Proposition 1 in García Rasines and Young (2022).
Proposition 3.
Let $Y\sim N(\theta,\sigma^{2})$, with $\sigma^{2}>0$ known, and $p(y)$ a non-constant, increasing, and right-continuous selection function with $\|p\|_{\infty}<\infty$. Let $\Pi(\theta\mid y)$ be the selective posterior distribution function based on the uniform prior $\pi(\theta)\propto 1$. Then,
$P_{\theta_{0}}\{\theta_{0}\leq\Pi^{-1}(\alpha\mid Y)\mid u\}<\alpha\quad\forall(\alpha,\theta_{0})\in(0,1)\times\mathbb{R}.$ (18)
The proof of this result hinges on the existence of a function $\pi(\theta;y)$ which is strictly increasing in $\theta$ for every $y$ such that, if employed as a data-dependent prior density, it achieves exact probability matching. That is, $P_{\theta_{0}}\{\theta_{0}\leq\Pi^{-1}(\alpha\mid Y)\mid u\}=\alpha$ for all possible values of $\theta$ and $\alpha$. Then, by comparison with the constant prior, it follows that the latter produces posterior inferences which are biased towards lower values of $\theta$ under repeated sampling.
If the selection function is increasing, naive selective inference based on the unconditional model $Y\sim N(\theta,\sigma^{2})$ would lead to overestimation of $\theta$. Proposition 3 shows that correcting for selection via conditioning on the selection event has the opposite effect in a Bayesian setting: it leads to systematic underestimation of $\theta$. Note that for a realized $Y=y$, $C(y)=(-\infty,\Pi^{-1}(\alpha\mid y)]$ is a $\alpha 100\%$ credible interval for $\theta_{0}$, so the proposition states that all intervals of this form undercover the parameter. In García Rasines and Young (2022) it is shown that certain default priors, which depend on the sample size and assign low prior probability to parameters with low selection probability, can be used to correct this effect.
In view of Corollary 1, a similar behavior ought to arise in more general parametric settings provided the induced selection function on the normalized score, $p^{*}_{n}(z;\theta)$, is increasing. Figure 5 illustrates this phenomenon in the exponential model considered above, where all the induced selection functions are increasing. Qualitatively, we observe that when selection is deterministic there is a large deviation between the distribution of $\Pi(\theta_{0}\mid Y)$ and the uniform distribution, which would correspond with a perfectly calibrated posterior. This behavior persists from the Gaussian model to the exponential model. Randomizing selection reduces this discrepancy, but the deviation remains significant.
Figure 5: Cumulative distribution functions of $\Pi(\theta_{0}\mid Y^{n})$ in the exponential model (blue) with the same settings as before, and of the corresponding Gaussian posterior distributions (black). Left, $n=50$; right, $n=100$.
5 Discussion
To date, Gaussianity assumptions have formed the basis of most analysis of inferential properties of selection models and the techniques of selective inference. The asymptotic analysis presented in this paper establishes formal connections between the behavior of selective inference methodologies in more general parametric settings and of their analogues in the corresponding Gaussian limits. The main focus for future work will be to utilize these connections to examine operational and theoretical optimality properties of procedures advocated for inference in the presence of selection. In particular, it will be of methodological importance to explore extensions of the main result to regression settings.
A key objective of the paper has been to provide insights on the methodological implications of selection for Bayesian inference. Specifically, the convenient asymptotic analogy between frequentist and Bayesian inference in regular random sampling models has been demonstrated to no longer hold under selection. This issue can be especially problematic in settings where the selection mechanism is consistently biased towards one region of the parameter space, as is often the case in practice under common selection conditions. The results presented in this paper point to the need for reexamination of conventional specifications of prior distributions, if appropriate frequentist calibration of the Bayesian inference, as is asymptotically achieved in regular IID problems without selection, is considered important.
Appendix A Appendix
A.1 Proof of Theorem 1
Fix a $\theta\in\Theta$ satisfying the required assumptions and let
$X_{n}(h)=\log\{f_{\theta+h/\sqrt{n}}(Y^{n})/f_{\theta}(Y^{n})\}.$
We have the unconditional expansion
$X_{n}(h)=h^{T}I_{\theta}Z_{n}-\frac{1}{2}h^{T}I_{\theta}h+o_{F_{\theta}}(1),$
where $Z_{n}=n^{-1/2}I_{\theta}^{-1}\sum_{i=1}^{n}\triangledown l_{\theta}(Y_{i})\rightsquigarrow N(0,I_{\theta}^{-1})$. We can write
$r_{n}(h)=\frac{\varphi_{n}(\theta+n^{-1/2}h)}{\varphi_{n}(\theta)}=E_{\theta}\left[\exp\left\{X_{n}(h)\right\}\mid u_{n}\right].$ (19)
Consider the sequence of random vectors $[Z,U^{*}_{n}]$, where $Z\sim N(h,I_{\theta}^{-1})$ and $U^{*}_{n}\mid Z\sim\text{Bernoulli}\{p^{*}_{n}(Z;\theta)\}$, so that $Z\mid U_{n}^{*}=1$ has selection function $p^{*}_{n}(z;\theta)$. The analogous functions for this model are given by
$X^{*}(h)$ $=$ $h^{T}I_{\theta}Z-\frac{1}{2}h^{T}I_{\theta}h;$ (20)
$r^{*}_{n}(h)$ $=$ $\frac{\varphi^{*}_{n}(h)}{\varphi^{*}_{n}(0)}=E_{0}\left[\exp\left\{X^{*}(h)\right\}\mid U^{*}_{n}=1\right].$ (21)
The main part of the proof concerns showing that $r_{n}(h_{n})-r^{*}_{n}(h_{n})\to 0$ for any convergent $h_{n}\to h$. To this end we show the following approximations:
(A) $E_{\theta}[\exp(h_{n}^{T}I_{\theta}Z_{n})p^{*}_{n}(Z_{n};\theta)]-E_{0}[\exp(h_{n}^{T}I_{\theta}Z)p^{*}_{n}(Z;\theta)]\to 0;$ (22)
(B) $E_{\theta}[(\exp\{X_{n}(h_{n})\}-\exp\{h_{n}^{T}I_{\theta}Z_{n}-(1/2)h_{n}^{T}I_{\theta}h_{n}\})p_{n}(Y^{n})]\to 0.$ (23)
(A) For all $K>0$, the collection $\{\exp(h^{T}Z_{n})\colon\|h\|\leq K,n\in\mathbb{N}\}$ is uniformly integrable under $\theta$. To see this, consider the moment generating function $M(x)=E_{\theta}[\exp\{x^{T}\triangledown l_{\theta}(Y_{1})\}]$. We have $E_{\theta}[\exp(h^{T}Z_{n})]=M(n^{-1/2}h)^{n}$, so it suffices to prove that
$\sup\{E_{\theta}[\exp(h^{T}Z_{n})^{2}]=M(n^{-1/2}2h)^{n}\colon|h|\leq K,n\in\mathbb{N}\}<\infty.$ (24)
Since $M(0)=1$ and $\triangledown M(0)=0$, a Taylor expansion of $\log M(x)$ gives
$n\log M\left(\frac{2h}{n^{1/2}}\right)=2h^{T}\left[\triangledown^{2}\log M(\tilde{x})\right]h,\quad\|\tilde{x}\|\leq\frac{2\|h\|}{n^{1/2}},$ (25)
so for some $K^{\prime}>0$ and any large enough $n$,
$\sup_{\|h\|\leq K,n\in\mathbb{N}}M\left(\frac{2h}{n^{1/2}}\right)^{n}\leq\exp\left\{2K^{2}\sup_{\|x\|\leq K^{\prime}}\triangledown^{2}\log M(x)\right\}<\infty.$ (26)
Clearly, $\{\exp(h^{T}I_{\theta}Z_{n})\colon\|h\|\leq K,n\in\mathbb{N}\}$ is also uniformly integrable for all $K>0$. Thus, for every $\varepsilon>0$, there exists a $C_{1}>0$ such that
$\sup_{\|h\|\leq K,n\in\mathbb{N}}E_{\theta}\left[\exp(h^{T}I_{\theta}Z_{n})\mathbf{1}\{\exp(h^{T}I_{\theta}Z_{n})>C_{1}\}\right]<\varepsilon.$ (27)
Similarly, there is a $C_{2}>0$ such that
$\sup_{\|h\|\leq K,n\in\mathbb{N}}E\left[\exp(h^{T}I_{\theta}Z)\mathbf{1}\{\exp(h^{T}I_{\theta}Z)>C_{2}\}\right]<\varepsilon.$ (28)
Let $C=\max\{C_{1},C_{2}\}$ and define
$g_{n}(z;h)=\exp(h^{T}I_{\theta}z)p^{*}_{n}(z;\theta)\mathbf{1}\{\exp(h^{T}I_{\theta}z)\leq C\},$ (29)
which satisfies $\|g_{n}(\cdot;h)\|_{\infty}\leq C\|p^{*}_{n}(\cdot;h)\|_{\infty}$. If the score has an absolutely continuous component, by Theorem 2.5 of Bally and Caramellino (2016), $d_{\text{TV}}(Z_{n},Z)\to 0$, so $E_{\theta}[g_{n}(Z;h)]-E_{\theta}[g_{n}(Z_{n};h)]=o(1)$. Otherwise, define $A(h)=\{z\colon\exp(h^{T}I_{\theta}z)\leq C\}$. We have that
$\|g_{n}(\cdot;h)_{|A(h)}\|_{\infty}$$\leq$ $C\|p^{*}_{n}(\cdot;\theta)\|_{\infty};$ (30)
$\|g_{n}(\cdot;h)_{|A(h)}\|_{\text{L}}$$\leq$ $C\left(\|p^{*}_{n}(\cdot;\theta)_{|A(h)}\|_{\text{L}}+\|I_{\theta}h\|\|p^{*}_{n}(\cdot;\theta)_{|A(h)}\|_{\infty}\right)$ (31)
$\leq$ $C\left(\|p^{*}_{n}(\cdot;\theta)\|_{\text{L}}+\|I_{\theta}h\|\|p^{*}_{n}(\cdot;\theta)\|_{\infty}\right);$ (32)
where we have used that $\|fg\|_{\text{L}}\leq\|f\|_{\infty}\|g\|_{\text{L}}+\|g\|_{\infty}\|f\|_{\text{L}}$. Since the bounded-Lipschitz distance metrizes convergence in distribution, we also get $E_{\theta}[g_{n}(Z;h)]-E_{\theta}[g_{n}(Z_{n};h)]=o(1)$. Thus, for all $\varepsilon>0$,
$\left|E_{\theta}[\exp(h_{n}Z_{n})p^{*}_{n}(Z_{n};\theta)]-E_{0}[\exp(h_{n}Z)p^{*}_{n}(Z;\theta)]\right|\leq 2\varepsilon\|p^{*}_{n}(\cdot;\theta)\|_{\infty}+o(1),$ (33)
so the term on the left hand side is $o(1)$, verifying (A).
In particular, for $h_{n}=0$, we get $\varphi_{n}(\theta)-\varphi_{n}^{*}(0)=o(1)$, which implies that $\varphi_{n}(\theta)/\varphi_{n}^{*}(0)=1+o(1)$ by the lower bound assumption on $\varphi_{n}(\theta)$.
(B) First, we have $E_{\theta}[\exp\{X_{n}(h)\}]=1$ for all $h$. Since $Z_{n}\rightsquigarrow Z$ and $\{\exp(h^{T}I_{\theta}Z_{n})\colon\|h\|\leq K,n\in\mathbb{N}\}$ is uniformly integrable, it follows that
$E_{\theta}[\exp\{h_{n}^{T}I_{\theta}Z_{n}-(1/2)h_{n}^{T}I_{\theta}h_{n}\}]\to E_{0}[\exp\{h^{T}I_{\theta}Z-(1/2)h^{T}I_{\theta}h\}]=1,$
so
$E_{\theta}[\exp\{X_{n}(h_{n})\}-\exp\{h_{n}^{T}I_{\theta}Z_{n}-(1/2)h_{n}^{T}I_{\theta}h_{n}\}]\to 0.$ (34)
Since $p_{n}(y^{n})$ is bounded, this implies the asserted claim.
Putting all together, using $\dot{=}$ to indicate equality up to an $o(1)$ term,
$\varphi_{n}(\theta)\{r_{n}(h_{n})-r_{n}^{*}(h_{n})\}$ (35)
$=$ $E_{\theta}[\exp\{X_{n}(h_{n})\}p_{n}(Y^{n})]-\frac{\varphi_{n}(\theta)}{\varphi^{*}_{n}(0)}E_{0}[\exp\{X^{*}(h_{n})\}p^{*}_{n}(Z;\theta)]$ (36)
$\dot{=}$ $E_{\theta}[\exp\{h_{n}^{T}I_{\theta}Z_{n}-(1/2)h_{n}^{T}I_{\theta}h_{n}\}p_{n}(Y^{n})]-E_{0}[\exp\{X^{*}(h_{n})\}p^{*}_{n}(Z;\theta)]$ (37)
$=$ $E_{\theta}[\exp\{h_{n}^{T}I_{\theta}Z_{n}-(1/2)h_{n}^{T}I_{\theta}h_{n}\}p^{*}_{n}(Z_{n};\theta)]-E_{0}[\exp\{X^{*}(h_{n})\}p^{*}_{n}(Z;\theta)]$ (38)
$\dot{=}$ $E_{\theta}[\exp\{h_{n}^{T}I_{\theta}Z_{n}\}p^{*}_{n}(Z_{n};\theta)]-E_{0}[\exp\{h_{n}^{T}I_{\theta}Z\}p^{*}_{n}(Z;\theta)]$ (39)
$=$ $o(1).$ (40)
Since $\varphi_{n}(\theta)$ is bounded away from zero, we obtain $r_{n}(h_{n})-r_{n}^{*}(h_{n})=o(1)$. This establishes the unconditional expansion
$\log\frac{f_{\theta+h_{n}/\sqrt{n}}(Y^{n}\mid u_{n})}{f_{\theta}(Y^{n}\mid u_{n})}=h^{T}I_{\theta}Z_{n}-\frac{1}{2}h^{T}I_{\theta}h+\log\frac{\varphi_{n}^{*}(0)}{\varphi_{n}^{*}(h)}+o_{F_{\theta}}(1).$ (41)
However, the lower boundedness assumption on $\varphi_{n}(\theta_{0})$ ensures that the remainder term is also $o_{F^{u_{n}}_{\theta}}(1)$.
Finally,
$\varphi_{n}(\theta)\{E_{\theta}[g(Z_{n})\mid u_{n}]-E_{0}[g(Z)\mid U_{n}^{*}=1]\}$ (42)
$=$ $E_{\theta}[g(Z_{n})p^{*}_{n}(Z_{n};\theta)]-\frac{\varphi_{n}(\theta)}{\varphi^{*}_{n}(0)}E_{0}[g(Z)p^{*}_{n}(Z;\theta)]$ (43)
vanishes uniformly for all bounded functions $g$ if Condition 1 is satisfied, showing the last assertion. Otherwise, the same statement holds for all bounded-Lipschitz functions $g$.
A.2 Proof of Lemma 1
Under DQM, the non-selective model is Locally Asymptotically Normal at $\theta$, so $f_{\theta}$ and $f_{\theta+h_{n}/\sqrt{n}}$ are contiguous for any bounded sequence $h_{n}$. By Le Cam’s first lemma, if $\varphi_{n}(\theta)=E_{\theta}[p_{n}(Y^{n})]=o(1)$, $\varphi_{n}(\theta+h_{n}/\sqrt{n})=E_{\theta+h_{n}/\sqrt{n}}[p_{n}(Y^{n})]=o(1)$ for any bounded sequence $h_{n}$. Let $H_{n}=n^{1/2}(\hat{\theta}_{n}-\theta)$. Then, $\varphi_{n}(\hat{\theta}_{n})=\varphi_{n}(\theta+H_{n}/\sqrt{n})$. By conditioning on events $\{\|H_{n}\|\leq M\}$ for an arbitrarily large $M$, it is easy to see that $\varphi_{n}(\theta)\to 0\Rightarrow\varphi_{n}(\hat{\theta}_{n})=o_{F^{u_{n}}_{\theta}}(1)$. Similarly, if $\varphi_{n}(\hat{\theta}_{n})=o_{F^{u_{n}}_{\theta}}(1)$, then there exists a bounded sequence $h_{n}$ for which $\varphi_{n}(\theta+h_{n}/\sqrt{n})=o(1)$, so $\varphi_{n}(\theta)=0$.
A.3 Proof of Proposition 1
The proof follows the arguments of van der Vaart (1998) (page 141) almost step by step, to which we refer the reader for the technical details. Here, we indicate the elements of the proof that require further justification. Let $\Pi_{n}$ denote the prior distribution for the local parameter $h=\sqrt{n}(\theta-\theta_{0})$. First, we need to show that asymptotically the posterior distributions of $h$ obtained with $\Pi_{n}$ are equivalent to those obtained with the restriction of $\Pi_{n}$ to $C_{n}$, the ball of radius $M_{n}$ around 0, for any $M_{n}\to\infty$. Then, we apply the local expansion given by Theorem 1.
For the first part, let $F^{u_{n}}_{n,h}$ denote the selective distribution of the data under $h$. Since $\varphi_{n}(\theta_{0})$ is bounded away from zero, it follows that $F^{u_{n}}_{n,h_{n}}\triangleleft\triangleright F^{u_{n}}_{n,0}$ for every bounded $h_{n}$. To see this, note that
$P_{\theta_{0}}(A_{n}\mid u_{n})\to 0$ (44)
$\iff$ $E_{\theta_{0}}[p_{n}(Y^{n})\mathbf{1}_{A_{n}}(Y^{n})]\to 0$ (45)
$\iff$ $E_{\theta_{0}+h/\sqrt{n}}[p_{n}(Y^{n})\mathbf{1}_{A_{n}}(Y^{n})]\to 0,$ (46)
as $F_{n,h_{n}}\triangleleft\triangleright F_{n,0}$. If $\varphi_{n}(\theta_{0}+h/\sqrt{n})\to 0$ then $\varphi_{n}(\theta_{0})\to 0$ by contiguity, so the last statement is equivalent to $P_{\theta_{0}+h/\sqrt{n}}(A_{n}\mid u_{n})\to 0$. Furthermore, we can extend Lemma 10.3 of van der Vaart (1998) to the sequence of selection models. Under the conditions of Corollary 1, for every $M_{n}\to\infty$, there exists a sequence of tests $T_{n}$ and a constant $c>0$ such that, for every sufficiently large $n$ and every $\sqrt{n}\|\theta-\theta_{0}\|\geq M_{n}$,
$E_{\theta_{0}}(T_{n})\to 0\quad\text{and}\quad E_{\theta}(1-T_{n})\leq e^{-cn(\|\theta-\theta_{0}\|^{2}\wedge 1)}.$ (47)
However, under the assumptions on the sequence of selection functions, this also holds conditionally on $u_{n}$, i.e.
$E_{\theta_{0}}(T_{n}\mid u_{n})\to 0\quad\text{and}\quad E_{\theta}(1-T_{n}\mid u_{n})\leq e^{-cn(\|\theta-\theta_{0}\|^{2}\wedge 1)}.$ (48)
Indeed, we have
$E_{\theta_{0}}(T_{n}\mid u_{n})=\frac{1}{\varphi(\theta_{0})}E_{\theta_{0}}[p_{n}(Y^{n})T_{n}]\to 0,$ (49)
as $p_{n}$ is uniformly bounded and $\varphi_{n}(\theta_{0})$ is bounded away from zero. Moreover, for some constant $K$,
$E_{\theta}(1-T_{n}\mid u_{n})=\frac{1}{\varphi(\theta_{0})}E_{\theta_{0}}[p_{n}(Y^{n})(1-T_{n})]\leq KE_{\theta}(1-T_{n}),$ (50)
so the second condition is also satisfied. This establishes the first claim.
To conclude the proof, let $C$ be a fixed ball around zero of fixed radius. Denote by $\Pi^{C}_{h,n}(\cdot\mid Y^{n})$ and $\Pi^{*C}_{n}(\cdot\mid Z_{n})$ the respective probability measures restricted to $C$, and use analogous notation for the corresponding densities. Their total variation distance can be bounded as
$\frac{1}{2}\left\|\Pi^{C}_{h,n}(\cdot\mid Y^{n})-\Pi^{*C}_{n}(\cdot\mid Z_{n})\right\|_{\mathrm{TV}}$
$\leq\int\int\left(1-\frac{\pi_{n}(g)f_{\theta_{0}+g/\sqrt{n}}(Y^{n}\mid u_{n})\pi^{*C}_{n}(h\mid Z_{n})}{\pi_{n}(h)f_{\theta_{0}+h/\sqrt{n}}(Y^{n}\mid u_{n})\pi^{*C}_{n}(g\mid Z_{n})}\right)^{+}\pi^{*C}_{n}(g\mid Z_{n})\pi^{C}_{h,n}(h\mid Y^{n})\mathrm{d}g\mathrm{d}h.$
By Theorem 1 and the regularity assumption on the prior, the integrand converges to zero in probability under the measure $\lambda_{C}(\mathrm{d}h)F_{\theta_{0}}^{u_{n}}(\mathrm{d}y^{n})\lambda_{C}(\mathrm{d}g)$, where $\lambda_{C}$ denotes the uniform measure on $C$. By the dominated convergence theorem, the right term of the inequality converges to zero in probability under the selective distribution, which concludes the proof, as this holds for an arbitrary $C>0$.
A.4 Proof of Proposition 2
For notational convenience, we are going to consider the two-dimensional case with
$\Sigma=\begin{pmatrix}\sigma_{1}^{2}&\rho\sigma_{1}\sigma_{2}\\ \rho\sigma_{1}\sigma_{2}&\sigma_{2}^{2}\end{pmatrix},$
and parameter of interest $\theta_{1}$, though all the steps are readily generalizable to $p>2$. First, let us transform the data to
$Z=\begin{pmatrix}1&0\\ -c\lambda&c\end{pmatrix}Y\sim N\left(\begin{pmatrix}\theta_{1}\\ c[\theta_{2}-\lambda\theta_{1}]\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right)\equiv N\left(\begin{pmatrix}\psi\\ \chi\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right),$
where
$\lambda$ $=$ $\rho\frac{\sigma_{2}}{\sigma_{1}},$
$c$ $=$ $\left(\sigma_{2}^{2}\left[1+\frac{\rho^{2}}{\sigma_{1}^{2}}(1-2\sigma_{1}^{2})\right]\right)^{-1/2},$
and $\psi=\theta_{1}$ and $\chi=c(\theta_{2}-\lambda\theta_{1})$ are the parameter of interest and the nuisance parameters, respectively. Let $\tilde{\varphi}(\psi,\chi)$ be the selection probability in the $(\psi,\chi)$ parametrization. We want to show that $\Pi(\psi_{0}\mid Z)$ is not distributed as $U(0,1)$ under any value of $\psi_{0}$.
The selective posterior density for $\psi$, for a given prior $\pi(\psi,\chi)$, has
$\pi(\psi\mid z_{1},z_{2})\propto\phi(\psi-z_{1})g(\psi;z_{2}),\quad g(\psi;z_{2})=\int_{-\infty}^{\infty}\pi(\psi,\chi)\frac{\phi(\chi-z_{2})}{\tilde{\varphi}(\psi,\chi)}\mathrm{d}\chi.$
We are going to show first that $\Pi(\psi_{0}\mid Z_{1},Z_{2})\mid Z_{2}=z_{2}\not\sim U(0,1)$ for some $z_{2}$ when $\psi_{0}$ is the true value of the interest parameter, irrespective of the value of $\psi_{0}$. The selective density of $Z_{1}$ given $Z_{2}=z_{2}$ is
$f_{\psi}(z_{1}\mid z_{2},u)=\frac{\phi(\psi-z_{1})\tilde{p}(z_{1},z_{2})}{\tilde{\varphi}(\psi;z_{2})},\quad\tilde{\varphi}(\psi;z_{2})=\int_{-\infty}^{\infty}\phi(\psi-z_{1})\tilde{p}(z_{1},z_{2})\mathrm{d}z_{1},$
where $\tilde{p}(z_{1},z_{2})$ is the selection function for the transformed data $Z$. Lindley (1958) showed that, for a one-dimensional model to admit an exact probability-matching prior, there needs to exist a change of variables $X=x(Z_{1})$ and a reparametrization $\tau=\tau(\psi)$ such that the model for $X\mid\tau$ is a location model. However, no transformation of $\psi$ can turn the conditional model of $Z_{1}\mid Z_{2}=z_{2}$ into a location model unless $\tilde{\varphi}(\psi;z_{2})\propto 1$, as the likelihood is not expressible in the form $L\{\tau-x^{-1}(X)\}$ for any $L$. Furthermore, since $\tilde{\varphi}(\psi,\chi)=E_{\chi}[\tilde{\varphi}(\psi;Z_{2})]$, we have, by the assumption on the selection probability, that
$E_{\chi}\left[\frac{\partial}{\partial\psi}\tilde{\varphi}(\psi;Z_{2})\right]$ $=$ $\frac{\partial}{\partial\psi}\tilde{\varphi}(\psi,\chi)$
$=$ $\frac{\partial}{\partial\theta_{1}}\varphi(\theta_{1},\theta_{2})\frac{\partial\theta_{1}}{\partial\psi}+\frac{\partial}{\partial\theta_{2}}\varphi(\theta_{1},\theta_{2})\frac{\partial\theta_{2}}{\partial\psi}$
$=$ $\frac{\partial}{\partial\theta_{1}}\varphi(\theta_{1},\theta_{2})+\lambda\frac{\partial}{\partial\theta_{2}}\varphi(\theta_{1},\theta_{2})$
$=$ $\frac{1}{\sigma_{1}^{2}}\left[\text{Cov}(Y_{1},Y_{1})\frac{\partial}{\partial\theta_{1}}\varphi(\theta_{1},\theta_{2})+\text{Cov}(Y_{1},Y_{2})\frac{\partial}{\partial\theta_{2}}\varphi(\theta_{1},\theta_{2})\right]$
$\neq$ 00 0
for all $(\psi,\chi)$. It follows that $\psi\to\tilde{\varphi}(\psi;Z_{2})$ cannot be constant in $\psi$ with probability one. If $z_{2}$ is such that $\tilde{\varphi}(\psi;Z_{2})$ is not constant, there does not exist any prior $\pi(\psi;z_{2})$ that provides exact frequentist matching when attached to the conditional likelihood of $Z_{1}\mid Z_{2}=z_{2}$. Since $\pi(\psi\mid z_{1},z_{2})$ can be obtained by combining the conditional likelihood with $\pi(\psi;z_{2})\propto\tilde{\varphi}(\psi;z_{2})g(\psi;z_{2})$, it follows that $\Pi(\psi_{0}\mid Z_{1},Z_{2})$ is not be uniform conditionally on $Z_{2}=z_{2}$, for a set of $z_{2}$’s of non-zero probability.
Finally, from the latter observation we show that $\Pi(\psi_{0}\mid Z_{1},Z_{2})$ cannot be uniform under all possible values of the nuisance parameter $\chi_{0}$, thus failing to satisfy the required matching condition. If $\Pi(\psi_{0}\mid Z_{1},Z_{2})\sim U(0,1)$, then
$P_{\psi_{0},\chi_{0}}\{\Pi(\psi_{0}\mid Z_{1},Z_{2})\leq\alpha\}=E_{\psi_{0},\chi_{0}}\left[P_{\psi_{0}}\{\Pi(\psi_{0}\mid Z_{1},Z_{2})\leq\alpha\mid Z_{2}\}\right]=\alpha,\quad\forall\alpha\in(0,1).$
However, $Z_{2}$ is complete for $\chi$, as it is the sufficient statistic of a natural exponential family. Hence, this equality can only hold if $P_{\psi_{0}}\{\Pi(\psi_{0}\mid Z_{1},Z_{2})\leq\alpha\mid Z_{2}\}=\alpha$ for all $\alpha$ with probability one, which contradicts the earlier statement.
A.5 Proof of Proposition 3
Assume for simplicity that $\sigma^{2}=1$. As in the proof of Theorem 1, consider the model for $(Y,U)$ where $U\mid Y\sim\text{Bernoulli}\{p(Y)\}$, and let $H(\theta;y)=P_{\theta}(Y\geq y\mid U=1)$. Note that $H(\theta;Y)\sim U(0,1)$ under $\theta$. Rewrite this as
$H(\theta;y)=\frac{\varphi(\theta;y)}{\varphi(\theta)},$ (51)
where $\varphi(\theta;y)=P_{\theta}(U=1,Y\geq y)$. Furthermore, define the function
$\pi(\theta;y)=-\frac{H_{\theta}(\theta;y)}{H_{y}(\theta;y)},$ (52)
where subscripts denote partial differentiation. Formally, $\pi$ can be thought of a data-dependent probability matching prior, in the sense that, when appended to the likelihood of $Y\mid U=1$, it produces, by construction, a posterior distribution function equal to $H(\theta;y)$, which satisfies the probability-matching condition, as $H(\theta;Y)\sim U(0,1)$. The proof boils down to showing that this probability-matching prior is increasing in $\theta$ for every fixed $y$.
The cumulant generating function of $Y\mid U=1$ is $K(t;\theta)=t^{2}/2+\theta t+\log\{\varphi(\theta+t)/\varphi(\theta)\}$, so in particular $\varphi_{\theta}(\theta)/\varphi(\theta)=E_{\theta}(Y\mid U=1)-\theta$. Analogously, $\varphi_{\theta}(\theta;y)/\varphi(\theta)=E_{\theta}(Y\mid U=1,Y\geq y)-\theta$, and $\phi(\theta-y)/\Phi(\theta-y)=E_{\theta}(Y\mid Y\geq y)-\theta$, where $\Phi$ and $\phi$ are the standard Gaussian CDF and PDF, respectively. Using these facts, a direct calculation gives
$\pi(\theta;y)\propto\frac{\varphi(\theta;y)}{\Phi(\theta-y)}\times\frac{E_{\theta}(Y\mid U=1,Y\geq y)-E_{\theta}(Y\mid U=1)}{E_{\theta}(Y\mid Y\geq y)-E_{\theta}(Y)},$ (53)
where the proportionality constant depends on $y$ but not on $\theta$. The first factor is increasing in $\theta$ for every $y$, as
$\frac{\partial}{\partial\theta}\log\frac{\varphi(\theta;y)}{\Phi(\theta-y)}=E_{\theta}(Y\mid U=1,Y\geq y)-E_{\theta}(Y\mid Y\geq y)>0.$ (54)
This holds because the selection function is increasing (and non-constant).
The second factor is also increasing. Assume without loss that $\|p\|_{\infty}=1$ (as the model is invariant to rescaling) and note that the selection condition can be restated as $Y\geq T$, where $T$ is independent of $Y$ and $P(T\leq t)=p(t)$. Write
$E_{\theta}(Y\mid Y\geq T)=E[E_{\theta}(Y\mid Y\geq T,T)]=\theta+E\left[g(\theta-T)\right],$ (55)
where $g(x)=\phi(x)/\Phi(x)$, and, similarly,
$E_{\theta}(Y\mid Y\geq T)=\theta+E\left[g(\theta-\max\{T,y\})\right].$ (56)
We therefore have that
$\frac{E_{\theta}(Y\mid U=1,Y\geq y)-E_{\theta}(Y\mid U=1)}{E_{\theta}(Y\mid Y\geq y)-E_{\theta}(Y)}$ (57)
$=$ $E\left[\frac{g(\theta-\max\{T,y\})-g(\theta-T)}{g(\theta-y)}\right]$ (58)
$=$ $p(y)E\left[1-\frac{g(\theta-T)}{g(\theta-y)}\mid T<y\right].$ (59)
The term inside the expectation is increasing in $\theta$ for every fixed $T<y$ and every $y$ (García Rasines and Young, 2022, Proposition 1), which concludes the proof.